OpenStudy (anonymous):

y value of 3sqrt(y^2)=(1/2y)?

OpenStudy (anonymous):

$$3\sqrt{y^2} = \frac{1}{2}y$$ is that your question?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

well sqrt(y^2) is just y correct?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

ok so 3y = 1/2 y

OpenStudy (anonymous):

that makes no sense

OpenStudy (alekos):

somethings wrong with the original equation show us a screen shot

OpenStudy (anonymous):

unless y is 0 i guess

OpenStudy (anonymous):

well im trying to find the y value

OpenStudy (anonymous):

ok then y can only equal 0 to make that equation true

OpenStudy (alekos):

I agree

OpenStudy (anonymous):

so the only possible way is to make y=0?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

alright tys

OpenStudy (anonymous):

yw